eviant:

The longer leg of a 30°-60°-90° triangle is 16sqrt3. How long is the shorter leg?

eviant:

\[16\sqrt{3}\]

eviant:

@dude @Vocaloid

dude:

If they gave you the long leg as \(16\sqrt{3}\) then the short leg is just the number in front of the sqrt3

jhonyy9:

 sin 30° = 1/2 so 1/2 = AB / 16sqrt3 AB = 16sqrt3/2 AB = 8sqrt3

jhonyy9:

@dude is this correct ?

jhonyy9:

@Vocaloid your opinion please about this ? thank you

dude:

Yeah, that’s much easier and I don’t think eviant has learned triangle trig

jhonyy9:

but - pardon @dude - than this is correct ,so this mean that what you ve wrote above not is - hope i see it right

dude:

Oh Whoops, uh they gave us a leg, not the hypotenuse